# Generative AI for Semantic Communication: Architecture, Challenges, and Outlook

### v2

@@ III-B Implementation Workflow
After introducing the functionalities of the three layers in GAI-SCN, we revisit its key rationales and showcase the workflow below step-by-step to guide network designers to make proper changes on relevant protocols, as shown in Fig. 4.
@@
Figure 4: A schematic diagram of implementing a complete round of semantic service provisioning in the GAI-SCN, including three successive stages of Initial Network Preparation, Generative AI-assisted SemCom Service Provisioning, and Model Synchronization and Update.
@@
Initial Network Preparation Stage: In this stage, the construction, pre-training and fine-tuning of GAI models are implemented collaboratively across the cloud and mobile layers. Note that all AI models are well pre-trained under society data public to users in the first place, like text from Wikipedia or books and images and videos from Instagram or YouTube, etc. Subsequently, considering the preference discrepancy between different users, personal data are collected by local GAI and then shared with global GAI in the cloud. Next, multiple parallel branch models corresponding to multiple users are created and fine-tuned based on personal data, where the branch model refers to a design pattern about multiple paths or subnetworks to make different predictions. Afterward, all parameters trained in each branch model are downloaded into the local light-weight GAI model related to each user, so as to avoid the drawback of limited computing ability in mobile TDs. Apart from the above, the construction and joint pre-training of semantic encoders and decoders are also completed in advance given specific channel state information. Keeping in mind the requirement of knowledge equivalence in SemCom, knowledge sharing between encoders and decoders is required if the condition is triggered [2]. GAI-Assisted SemCom Service Provisioning Stage: Once all communication parties in the GAI-SCN are equipped with mature-trained AI models, the SemCom transmission services begin. The local GAI model first extracts keywords and identifies the explicit goal (referring to the exemplification in Fig. 3) for each user. Hereafter, such keywords and goals are uploaded via traditional communications to global GAI models in the cloud layer as input prompts, so that the corresponding branch AI model can create AIGC in line with the user preference to recover the original meaning with the original modality. When it comes to the downlink side, each edge server detects the current wireless channel state and analyzes the received AIGC service-related signaling to select an appropriate pre-trained semantic encoder. As such, the entire AIGC is smoothly encoded into semantic features to be transmitted to the corresponding TD. Since semantic errors may still occur in the above JSCC process due to potential signal impairment and the limited computing power of TD, under the assistance of personalized knowledge and user preference, its local generative-AI model is utilized for further semantic calibration, such as error correction for text content, color rendering for image content, and frame completion for video content, to enhance the resilience and robustness against semantic ambiguity for SemCom. Model Synchronization and Update Stage: At the end of each GAI-assisted SemCom process, the service experience feedback (i.e., the user satisfaction regarding service performance, availability, and accessibility, etc.) is collected from each user and cached in the associated edge server. The feedback data as well as real-time user transmission data are synced periodically from the edge layer to the GAI models deployed in the other two layers. In this way, GAI analyzes and predicts users’ up-to-date preferences by fine-tuning model parameters and structures. Meanwhile, the data are also fed back to semantic models to update the relative parameters as well as the linked background knowledge.
@@ IV Case Study: Image Transmission Service Provisioning in GAI-SCN
In this section, numerical results for a case study of image transmission are presented to evaluate the initial performance of the proposed GAI-SCN framework. For the simulation settings, an image-captioning model by combining ViT model with GPT-2 model [13] is exploited as the local GAI to realize the image-to-text transformation as well as the keyword extraction and goal identification. Besides, we employ the latest text-to-image model called Stable Diffusion 2.1 [6] as the global GAI to create the AI-generated images from received prompts. As for the SemCom part, the main setups are proceeding as in the work [2], where an advanced deep convolutional network named Observation-Centric Sort and a Transformer-powered semantic decoder are leveraged for semantic segmentation and recovery, respectively. Meanwhile, all semantic models are trained based on the additive white Gaussian noise channel with a signal-to-noise ratio of 00 dB to transmit $327$ images with different contents for testing. Finally, the Adam optimizer is adopted to train the neural networks in GAI-SCN with an initial learning rate of $5\times 10^{-4}$ based on the given image dataset. In parallel, for comparison purposes, we utilize two benchmarks: 1) A GAI-assisted traditional communication scheme, where the AIGC should be encoded into bits based on the prescribed coding rule [14] for precise image delivery; 2) A typical SemCom scheme [1], which transmits the original images solely via semantic coding models without any participation of GAI.
@@
Figure 5 first demonstrates the performance of GAI-SCN by comparing the recovered images with the original ones under different numbers of observable objects that can be detected and segmented (such as “bear”, “tree”, and “ground” as shown in the exemplification image in Fig. 3). Note that the appearances of the transmitted images may be not completely consistent with the received ones, however, the semantics implicit in the content should be our sole focus in measuring the system performance. Herein, we first test the semantic similarity performance measured by spaCy [15], where generally, the higher the spaCy score, the more accurate the recovered semantics. It is observed that the increasing number of objects in original images results in a decreasing semantic similarity, which is because that the higher complexity of images makes GAI more difficult to extract the keywords correctly as well as the image recovery. This phenomenon is also consistent with the recovery ratio performance of original objects, which metric shows the proportion of objects accurately recovered in the AIGC, and as the complexity of images rises, the average recovery ratio drops from $82.2\%$ to $41.8\%$ steadily. Moreover, we can see a lower object quantity discrepancy with fewer number of objects contained in the transmitted images. All of these trends above represent that semantic ambiguity is more likely to occur in regenerating more complex images due to confusing key object identification and a certain degree of semantic interference.
@@
GAI-assisted traditional communication [14]
$1.28\times 10^{5}$
$28.05$
SemCom [1]
$5.99\times 10^{4}$
$28.25$
GAI-SCN
$3.03\times 10^{4}$
$28.64$
In addition, in order to validate the superiority of integrating SemCom with GAI, Table I shows the number of transmission bits required for three different schemes on the downlink side as well as the corresponding peak signal-to-noise ratio (PSNR) performance. Notably, all results in this table are based on the test of transmitting $300$ images with a given size of $1024*1024$ pixels. It can be found that the proposed GAI-SCN only requires $3.03\times 10^{4}$ bits, which reduces $2.96\times 10^{4}$ bits compared with the SemCom scheme and $9.77\times 10^{4}$ bits with the GAI-assisted conventional scheme. Furthermore, the PSNR score obtained by our GAI-SCN maintains a very high level of $28.64$, which is even slightly better than the other two approaches. This can be explained by the fact that the typical SemCom scheme starts the image delivery process from the uplink direction, which increases the risk of image sharpness loss, while the GAI-assisted traditional communication scheme is less resilient to harsh channel conditions compared with GAI-SCN. In summary, the above results demonstrate that our GAI-SCN can further save bandwidth resources while guaranteeing very high-quality SemCom service provisioning with accurate semantic delivery.

### v6

@@ III-B Implementation Workflow
By revisiting the key rationales of GAI-SCN, we now showcase its workflow below step-by-step to guide network designers to make proper changes on relevant protocols, as shown in Fig. 4.
@@
Figure 4: A schematic diagram of implementing a complete round of semantic service provisioning in the GAI-SCN, including three successive stages of Initial Network Preparation, Generative AI-integrated SemCom Service Provisioning, and Model Synchronization and Update.
@@
Initial Network Preparation Stage: In this stage, the construction, pre-training and fine-tuning of GAI models are implemented collaboratively across the cloud and mobile layers. Note that all AI models are well pre-trained under society data public to users in the first place. Subsequently, considering the preference discrepancy between different users, personal data are collected by local GAI and then shared with global GAI in the cloud. Next, multiple parallel branch models corresponding to multiple users are created and fine-tuned based on personal data, where the branch model refers to a design pattern about multiple paths or subnetworks to make different predictions. Afterward, all parameters trained in each branch model are downloaded into the local light-weight GAI model related to each user, so as to avoid the drawback of limited computing ability in mobile TDs. Apart from the above, the construction and joint pre-training of semantic encoders and decoders are also completed in advance given specific channel state information. Keeping in mind the requirement of knowledge equivalence in SemCom, knowledge sharing between encoders and decoders is required if the condition is triggered [7].
GAI-Integrated SemCom Service Provisioning Stage: Once all communication parties in the GAI-SCN are equipped with mature-trained AI models, the SemCom transmission services begin. The local GAI model first extracts keywords and identifies the explicit goal (referring to the exemplification in Fig. 3) for each user. Hereafter, such keywords and goals are uploaded via traditional communications to global GAI models in the cloud layer as input prompts, so that the corresponding branch AI model can create AIGC in line with the user preference to recover the original meaning with the original modality.
When it comes to the downlink side, each edge server detects the current wireless channel state and analyzes the received AIGC service-related signaling to select an appropriate pre-trained semantic encoder. As such, the entire AIGC is smoothly encoded into semantic features to be transmitted to the corresponding TD. Since semantic errors may still occur in the above JSCC process due to potential signal impairment and the limited computing power of TD, under the assistance of personalized knowledge and user preference, its local generative-AI model is utilized for further semantic calibration, such as error correction for text content, color rendering for image content, and frame completion for video content, to enhance the resilience and robustness against semantic ambiguity for SemCom.
Model Synchronization and Update Stage: At the end of each GAI-integrated SemCom process, the service experience feedback (i.e., the user satisfaction regarding service performance, availability, and accessibility, etc.) is collected from each user and cached in the associated edge server. The feedback data as well as real-time user transmission data are synced periodically from the edge layer to the GAI models deployed in the other two layers. In this way, GAI analyzes and predicts users’ up-to-date preferences by fine-tuning model parameters and structures. Meanwhile, the data are also fed back to semantic models to update the relative parameters as well as the linked background knowledge.
@@ IV Case Study: Image Transmission Service Provisioning in GAI-SCN
In this section, numerical results for a case study of image transmission are presented to evaluate the initial performance of the proposed GAI-SCN framework. For the simulation settings, an image-captioning model by combining ViT model with GPT-2 model [13] is exploited as the local GAI to realize the image-to-text transformation as well as the keyword extraction and goal identification. Besides, we employ the latest text-to-image model called Stable Diffusion 2.1 [5] as the global GAI to create the AI-generated images from received prompts. As for the SemCom part, the main setups are proceeding as in the work [7], where an advanced deep convolutional network named Observation-Centric Sort and a Transformer-powered semantic decoder are leveraged for semantic segmentation and recovery, respectively. Meanwhile, all semantic models are trained based on the additive white Gaussian noise channel with a signal-to-noise ratio (SNR) of 00 dB to transmit $327$ images with different contents for testing. Finally, the Adam optimizer is adopted to train the neural networks in GAI-SCN with an initial learning rate of $5\times 10^{-4}$ based on the given image dataset. In parallel, for comparison purposes, we utilize two benchmarks: 1) A GAI-integrated traditional communication scheme, where the AIGC is encoded into bits based on the variable length source coding and LDPC channel coding [14] for precise image delivery; 2) A typical SemCom scheme [1, 3], which transmits the original images via only DL-based semantic coding and JSCC without any involvement of GAI.
@@
Figure 5 first demonstrates the performance of GAI-SCN by comparing the recovered images with the original ones under different numbers of observable objects that can be detected and segmented (such as “bear”, “tree”, and “ground” as shown in the exemplification image in Fig. 3). Note that the appearances of the transmitted images may be not completely consistent with the received ones, however, the semantics implicit in the content should be our sole focus in measuring the system performance. Herein, we first test the semantic similarity performance measured by spaCy [15], where generally, the higher the spaCy score, the more accurate the recovered semantics. Due to the cross-layer design of our proposed GAI-SCN framework, semantic similarity is actually an application-layer indicator to measure the degree of user semantic provisioning, which should be distinguished from lower-level technical indicators such as bit/symbol error ratios. Particularly, the error occurs at the bit/symbol level does not necessarily imply an error at the semantic level, while there could be errors at the semantic level even if there is no error at the bit level. The former is owing to the powerful semantic reasoning ability of semantic models and the assistance of perfectly matching background knowledge, and the latter is because of the potential background knowledge discrepancy between the sender and the receiver. Back to Fig. 5, it is observed that the increasing number of objects in original images results in a decreasing semantic similarity, which is because that the higher complexity of images makes GAI more difficult to extract the keywords correctly as well as the image recovery. This phenomenon is also consistent with the recovery ratio performance of original objects, which metric shows the proportion of objects accurately recovered in the AIGC, and as the complexity of images rises, the average recovery ratio drops from $82.2\%$ to $41.8\%$ steadily. Moreover, we can see a lower object quantity discrepancy with fewer number of objects contained in the transmitted images. All of these trends above represent that semantic ambiguity is more likely to occur in regenerating more complex images due to confusing key object identification and a certain degree of semantic interference.
@@
GAI-integrated traditional communication [14] $1.28\times 10^{5}$ $28.05$
SemCom [1] $5.99\times 10^{4}$ $28.25$
GAI-SCN $3.03\times 10^{4}$ $28.64$
Besides the above AI-generated quality measurement, the proposed GAI-SCN is further tested from the wireless transmission perspective. Table I shows the number of bits required and the image transmission performance achieved to downlink $300$ images with the equal size of $1024*1024$ pixels in the same 00 dB SNR Gaussian noise channel. For tractability, the peak signal-to-noise ratio (PSNR) is employed here to compare the images before and after performing different transmission schemes, noting that other image quality evaluation metrics like structural similarity index can also be adopted. Moreover, since we concentrate only upon the discrepancy in transmitted content (AIGC vs. original) and communication technology used (SemCom vs. traditional), the same source and channel coding rules are assumed for all schemes to encode the images from the pixel/semantic level to the bit level. It can be found that the proposed GAI-SCN only requires $3.03\times 10^{4}$ bits, which reduces $2.96\times 10^{4}$ bits compared with the SemCom scheme and $9.77\times 10^{4}$ bits with the GAI-integrated conventional scheme. Furthermore, the PSNR score obtained by our GAI-SCN maintains a very high level of $28.64$, which is even slightly better than the other two approaches. This can be explained by the fact that the typical SemCom scheme starts the image delivery process from the uplink direction, which increases the risk of image sharpness loss, while the GAI-integrated traditional communication scheme is less resilient to harsh channel conditions compared with GAI-SCN. In summary, the above results demonstrate that our GAI-SCN can further save bandwidth resources while guaranteeing very high-quality SemCom service provisioning with accurate semantic delivery.
@@ VI Conclusions
VII Acknowledgments
This work was supported in part by UK Department for Science, Innovation & Technology (DSIT) Towards Ubiquitous 3D Open Resilient Network (TUDOR) Project, and in part by the National Research Foundation, Singapore, and Infocomm Media Development Authority under its Future Communications Research & Development Programme, Defence Science Organisation (DSO) National Laboratories under the AI Singapore Programme (FCP-NTU-RG-2022-010 and FCP-ASTAR-TG-2022-003), Singapore Ministry of Education (MOE) Tier 1 (RG87/22), and the NTU Centre for Computational Technologies in Finance (NTU-CCTF).